OpenStudy (anonymous):

E= hf-w W= ? options a) E/hf B) -E/hf c) E+hf D) hf-E choose the right option from above.

OpenStudy (anonymous):

Surely you can subtract by yourself...

OpenStudy (anonymous):

i just dont understand the concept

OpenStudy (anonymous):

A = B - C so C = B - A ?

OpenStudy (anonymous):

so its the first option?

OpenStudy (anonymous):

No.... it's literally just addition and subtraction, there's no division involved...

OpenStudy (anonymous):

okay so then it would be the last one?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thankyou . would it be alright if i post a few more questions. ill tell u the answer plz just guide me?

OpenStudy (anonymous):

I don't need the answers. If you post them, either I or somebody else will get to them.

OpenStudy (anonymous):

Thanks